Find the least squares solution to the inconsistent system Ax = b 1 1 0 110 3 A = b = 101 8 1101 2
Answer #1

A 338 10) L101 Ax=b & ATAN=Atb 4 2 3 2 2 L 2 02 4 10 * 0 5 1-1/ 2-3 0 O [B[c] = wef (6/6) 0 o the = 5-43 Br = - 32 least sq.